A ball kicked horizontally at 10.0m/s from a cliff 80m high. How far from the base of the cliff will the stone strike the ground?


1
Expert's answer
2022-04-14T17:00:56-0400

"h(t)=H-\frac{gt^2}{2}=80-4.9t^2"

"h(t)=0 \to 80-4.9t^2 \to t=\sqrt{\frac{80}{4.9}}=4.04 s."

"S=vt=10*4.04=40.4m."